# FALCON: Resolving Visual Redundancy and Fragmentation in High-resolution Multimodal Large Language Models via Visual Registers

### v1

@@ Abstract
The incorporation of high-resolution visual input equips multimodal large language models (MLLMs) with enhanced visual perception capabilities for real-world tasks. However, most existing high-resolution MLLMs rely on a cropping-based approach to process images, which leads to fragmented visual encoding and a sharp increase in redundant tokens. To tackle these issues, we propose the FALCON model. FALCON introduces a novel visual register technique to simultaneously: 1) Eliminate redundant tokens at the stage of visual encoding. To directly address the visual redundancy present in the output of vision encoder, we propose a Register-based Representation Compacting (ReCompact) mechanism. This mechanism introduces a set of learnable visual registers designed to adaptively aggregate essential information while discarding redundancy. It enables the encoder to produce a more compact visual representation with a minimal number of output tokens, thus eliminating the need for an additional compression module. 2) Ensure continuity in visual encoding. To address the potential encoding errors caused by fragmented visual inputs, we develop a Register Interactive Attention (ReAtten) module. This module facilitates effective and efficient information exchange across sub-images by enabling interactions between visual registers. It ensures the continuity of visual semantics throughout the encoding. We conduct comprehensive experiments with FALCON on high-resolution benchmarks across a wide range of scenarios. FALCON demonstrates superior performance with a remarkable 9-fold and 16-fold reduction in visual tokens.
@@ 1 Introduction
Recently, multimodal large language models (MLLMs) [31, 29, 3, 64, 65, 8] have made significant advancements in performing vision-language tasks, facilitating more natural human-machine interactions. Early MLLMs [31, 64, 8] typically processed images at a fixed low resolution, such as 224×224. This limited their capabilities in many real-world applications that require high-resolution input. To equip MLLMs with high-resolution visual perception, recent advancements [32, 63, 14, 69, 17] have introduced a cropping-based high-resolution processing method. Specifically, As shown in Fig. 1(b), this approach crops high-resolution images into multiple non-overlapping sub-images that match the pre-trained resolution of the encoder. Each sub-image is processed independently by the vision encoder. Then the resulting visual features are concatenated and fed into the large language model (LLM).
@@
  • To address the issue of visual redundancy directly at the encoding phase, we introduce the Register-base Representation Compacting (ReCompact) mechanism. This mechanism enables vision encoder to capture rich visual information using only a minimal number of tokens, effectively eliminating redundancy and obviating the need for additional compression modules.
@@
  • We introduce FALCON, a model that leverages the proposed visual register technique. Comprehensive evaluations on high-resolution visual understanding benchmarks highlight the superiority of FALCON. Notably, FALCON achieves exceptional results with a remarkable reduction in visual tokens by 9× and 16×.
@@ 2.1 Multimodal Large Language Models
Research on MLLMs [29, 3, 2, 64, 65, 47] has made significant strides in recent years. By incorporating image inputs into pretrained LLMs [55, 56, 7], MLLMs demonstrate strong capabilities across various vision-language tasks. Typically, these models use a pretrained vision encoder [45, 66] to transform input images into a set of visual tokens. These tokens are then projected into the input space of the LLM through a projection module (e.g., MLP [31, 29], QFormer [8] and Abstractor [65]) and processed alongside text inputs. Building on the progress of MLLMs, recent studies have started to explore their applications in practical tasks such as document understanding [62, 14], agents [27], and autonomous driving [59]. Despite these advances, MLLMs still struggle with the fixed low resolution of pretrained image encoders, which can limit their effectiveness in handling high-resolution images.
@@ 3.1 Mitigating Visual Redundancy via ReCompact
Figure 3: The progressive training pipeline. Stage 1: Pretrain visual registers and MLP on low-resolution for coarse alignment. Stage 2: Tune all parameters with caption and OCR data on low-resolution for refined alignment. Stage 3: Integrate high-resolution input and ReAtten module for instruction tuning.
@@
The proposed ReCompact mechanism differs significantly from widely used Abstractor-like modules [65, 63, 2, 32, 28, 8]. While both share the goal of generating compact and non-redundant visual representations, the ReCompact shows notable advantages. Abstractor-like modules typically consist of multiple stacked Transformer [57] blocks, serving as an additional compression module between the vision encoder and the LLM. These modules use a set of query tokens to extract information from output image tokens through cross-attention in each block. In contrast, based on the observation that a well-pretrained ViT can naturally gather global information into some specific tokens through self-attention [9], ReCompact directly adopts the pretrained ViT to aggregate information into a set of specified register tokens. This design offers ReCompact several strengths in terms of: 1) Parameter Efficiency. Introducing abstractor-like modules significantly increases the number of parameters, whereas ReCompact avoids this issue by directly adapting the pretrained ViT. This mitigates the hardware burden in terms of both time and space. 2) Training Efficiency. The substantial number of randomly initialized parameters in Abstractor-like modules requires extensive data for pretraining. For instance, Qformer [23], Abstractor [65], and Resampler [2] use large corpora of 129M, 400M, and 1.4B samples, respectively. In contrast, by leveraging well-pretrained ViT parameters, ReCompact requires fewer than 2M samples for adaptation. The ablation results in Fig. 4 further demonstrate these advantages.
@@ 3.3 Learning Expressive Visual Representation via Progressive Training Pipeline
To empower the ReCompact mechanism with the capability to capture rich visual patterns and to enable smoother initialization of the ReAtten module under high-resolution conditions, we develop a specialized progressive training pipeline. As illustrated in Fig. 3, this pipeline progressively transitions from low to high resolution across three stages.
Stage 1: Coarse Alignment on Low-resolution. Since both the visual registers and the MLP module are randomly initialized, in stage 1, we aim to initially achieve a coarse alignment between the pretrained vision encoder and the LLM. To facilitate this, we begin by using low-resolution images as inputs, which are directly resized to match the pretraining resolution of the vision encoder. We then train the visual registers and the MLP using image caption data, while keeping both the vision encoder and the LLM frozen.
Stage 2: Refined Alignment on Low-resolution. To better adapt the pre-trained vision encoder to our newly introduced ReCompact mechanism, in stage 2, we unfreeze all model parameters to enable a more refined alignment. To enhance the capacity of the visual register to capture detailed and comprehensive visual information, we employ high-quality, extended captions that are rich in fine-grained details, along with a diverse set of full-document OCR data. Additionally, to ensure resource efficiency, we continue to use low-resolution image inputs. By combining these low-resolution inputs with the compact visual representations produced by the ReCompact mechanism, this approach facilitates efficient full parameter fine-tuning.
Stage 3: Instruction Tuning on High-resolution. To adapt the well-aligned visual register for a wide range of real-world vision-language tasks, in stage 3, we transition the model to process high-resolution image inputs. In this stage, we employ the cropping-based approach to empower the model with enhanced capability of high-resolution perception. Additionally, the proposed ReAtten module is integrated to address visual fragmentation caused by cropping-based high-resolution image encoding. To improve instruction-following performance, the training process utilizes a mixture of instruction-tuning datasets covering various tasks, such as VQA, document understanding, and grounding. To maintain resource efficiency during high-resolution training, the parameters of the vision encoder and the LLM are kept frozen, while only the LoRA parameters of the LLM and the remaining parameters are trainable. Notably, to facilitate a smoother initialization for the newly introduced ReAtten modules, we initialize parameters of each ReAtten module with the ViT self-attention parameters at the corresponding depth.
@@ 4 Experiments
We implement FALCON on pretrained LLMs of two different sizes, resulting in FALCON-7B and FALCON-3B. The numbers of visual registers for FALCON-7B and FALCON-3B are set to 64 and 36, respectively, achieving remarkable visual token compression rates of 9×9$\times$9 × and 16×16$\times$16 ×. For detailed information regarding the architecture, datasets and training settings, please refer to the Appendix.
@@ 4.1 Experimental Results
Table 1: Evaluation on MME-RealWorld Benchmark. “OCR”, “RS”, “DT”, “MO”, and “AD” stand for OCR in the Wild, Remote Sensing, Diagram and Table, Monitoring, and Autonomous Driving, respectively. “Avg” denotes the weighted average accuracy across all samples, while “Avg-C” denotes the weighted average accuracy across different domains. “$\text{Size}^{v}$” indicates the image resolution supported by the model, where ×$\times n$ signifies a cropping-based high-resolution processing approach with up to $n$ sub-images. “$\text{Token}^{c}$” denotes the number of visual tokens after compression. † For models utilizing hybrid vision encoders, “$\text{Size}^{v}$” and “$\text{Token}^{c}$” represent the setting of encoder with most tokens. We gray out the models that are either closed-source or trained on a corpus that is at least three times larger than ours. Among these, lighter gray denotes scores below ours. The underline indicates that top-3 performance is achieved. Notably, FALCON-7B achieves state-of-the-art performance with just 64 tokens per sub-image, surpassing even some advanced closed-source commercial models. Meanwhile, FALCON-3B also demonstrates impressive results, achieving a ranking just below FALCON-7B, despite having significantly fewer parameters and visual tokens—only 36 tokens per sub-image.
@@
Training Data $<$ 5M MiniGPT-v2 [5] $448^{2}$ $256$ 39.0 23.3 20.4 19.3 26.0 26.9 25.6 30.0 20.4 16.9 23.7 23.0 22.7
@@
FALCON-3B $336^{2}\times 16$ $\textbf{36}\times 16$ 57.3 35.7 34.5 31.6 33.1 40.5 38.4 48.2 29.8 19.9 31.7 32.2 32.4
FALCON-7B $336^{2}\times 16$ $\textbf{64}\times 16$ 60.1 43.5 41.4 34.4 39.5 45.9 43.8 50.8 29.8 32.9 31.7 34.8 36.2
Training Data $\gg$10M Qwen-VL-Chat [2] $448^{2}$ $256$ 32.4 15.1 15.6 22.1 15.1 20.8 20.1 28.6 13.6 16.5 24.6 22.0 20.8
Cambrian-1-8B† [54] $1024^{2}$ $576$ 58.7 40.1 32.7 47.7 38.5 43.8 43.5 53.2 27.4 42.4 30.7 36.2 38.4
@@
As shown in Tab. 1, FALCON-7B and FALCON-3B significantly outperform other advanced MLLMs in both perception and reasoning tasks. FALCON-7B achieves state-of-the-art performance with only 64 tokens per sub-image, resulting in a remarkable visual compression rate of 9 times. Furthermore, FALCON-3B achieves the second-best performance while using significantly fewer parameters of 3B and an even more impressive count of just 36 tokens per sub-image. In addition, even compared to models trained on much larger datasets, FALCON remains highly competitive. Specifically, FALCON-7B outperforms CogVLM2-llama3-Chat [13] in average scores on the perception task, which uses over 1 billion samples for training. It also surpasses MiniCPM-V 2.5 [61] and IXC2.5 [67] in average scores on the reasoning tasks, which utilize hundreds of millions of samples. FALCON-7B also exceeds all models trained on large corpora in several sub-tasks, including remote sensing and autonomous driving. Notably, FALCON achieves higher average scores in both perception and reasoning tasks compared to two advanced commercial models, GPT-4o-mini and Gemini-1.5-pro [53], with FALCON-7B notably outperforming all commercial models in the sub-tasks of remote sensing and monitoring. These results demonstrate that the ReCompact and ReAtten technique proposed by FALCON effectively eliminate visual redundancy and ensure visual continuity, allowing comprehensive and detailed visual information to be conveyed with a highly compact visual representation.
To further assess FALCON’s capabilities in high-resolution understanding, we conduct evaluations on three additional high-resolution benchmarks: V* [58], DocVQA [41], and TextVQA [50]. V* encompasses tasks related to attribute recognition and spatial reasoning in natural scenes, focusing on evaluating the models’ capabilities for detailed visual analysis. DocVQA challenges models to comprehend text-dense, high-resolution document images, while TextVQA evaluates the ability to recognize small text in real-world scenes. As shown in Tab. 2, FALCON demonstrates superior performance across all these benchmarks.
To evaluate the performance of FALCON across a broader range of domains, we further select three popular benchmarks: POPE [25], ScienceQA [37], and MMBench [33]. These benchmarks specifically assess models’ capabilities in hallucination, scientific reasoning, and comprehensive ability, respectively. As shown in Tab. 3, FALCON-7B and FALCON-3B achieve state-of-the-art and second-best performance across the three benchmarks, respectively. These results demonstrate that FALCON excels not only in high-resolution understanding but also exhibits superior capabilities across a wide range of tasks.
Table 2: Evaluations on diverse high-resolution Benchmarks. “$\textbf{V}^{*}_{Avg}$” represents the weighted average score for the two sub-tasks in the $V^{*}$ benchmark. * The TextVQA score of Monkey [28] is sourced from [16]. † denotes OCR inputs are utilized.
Model $\textbf{V}^{*}_{Avg}$ DocVQA TextVQA
@@
Qwen-VL [2] - 65.1 63.8
InstructBLIP [8] 34.0 4.5 50.7
LLaVA-v1.5-7B [31] 48.7 8.5 58.2†
mPLUG-Owl2 [65] - 17.9 58.2†
@@
UReader [63] - 65.4 57.6
Monkey [28] - 66.5 64.3*
LlaVA-FlexAttn [24] 54.4 - 48.9
S2-Wrapper-7B [48] 55.5 - 61.0
LlaVA-Next-Vicuna7B [30] 61.8 - 64.9
SliME-7B [69] - - 64.4
FALCON-3B 55.0 68.0 64.1
FALCON-7B 63.3 67.9 65.8

### v2

@@ Abstract
The incorporation of high-resolution visual input equips multimodal large language models (MLLMs) with enhanced visual perception capabilities for real-world tasks. However, most existing high-resolution MLLMs rely on a cropping-based approach to process images, which leads to fragmented visual encoding and a sharp increase in redundant tokens. To tackle these issues, we propose the FALCON model. FALCON introduces a novel visual register technique to simultaneously: 1) Eliminate redundant tokens at the stage of visual encoding. To directly address the visual redundancy present in the output of vision encoder, we propose a Register-based Representation Compacting (ReCompact) mechanism. This mechanism introduces a set of learnable visual registers designed to adaptively aggregate essential information while discarding redundancy. It enables the encoder to produce a more compact visual representation with a minimal number of output tokens, thus eliminating the need for an additional compression module. 2) Ensure continuity in visual encoding. To address the potential encoding errors caused by fragmented visual inputs, we develop a Register Interactive Attention (ReAtten) module. This module facilitates effective and efficient information exchange across sub-images by enabling interactions between visual registers. It ensures the continuity of visual semantics throughout the encoding. We conduct comprehensive experiments with FALCON on high-resolution benchmarks across a wide range of scenarios. FALCON demonstrates superior performance with a remarkable 9-fold reduction in visual tokens.
@@ 1 Introduction
Recently, multimodal large language models (MLLMs) [32, 30, 2, 58, 59, 9] have made significant advancements in performing various vision-language tasks [44, 45, 60], facilitating more natural human-machine interactions. Early MLLMs [32, 58, 9] typically processed images at a fixed low resolution, such as 224×224. This limited their capabilities in many real-world applications that require high-resolution input. To equip MLLMs with high-resolution visual perception, recent efforts [33, 57, 15, 66, 17] have introduced a cropping-based high-resolution processing method. Specifically, As shown in Fig. 1(b), this approach crops high-resolution images into multiple non-overlapping sub-images that match the pre-trained resolution of the encoder. Each sub-image is processed independently by the vision encoder. Then the resulting visual features are concatenated and fed into the large language model (LLM).
@@
Note that the widely used Abstractor-like modules [59, 57, 33, 28], including QFormer [9], ReSampler [1], follow a similar approach by using query tokens to extract information from image tokens after visual encoding. The key difference is that the visual register mechanism directly adopts the pretrained ViT to aggregate information into a set of specified register tokens. This design is based on the observation that a well-pretrained ViT can naturally gather global information into some specific tokens through self-attention [10]. It offers the visual register mechanism several strengths in terms of: 1) Training Efficiency. The substantial number of randomly initialized parameters in Abstractor-like modules requires extensive data for pretraining. For instance, QFormer [22], Abstractor [59], and Resampler [1] use large corpora of 129M, 400M, and 1.4B samples, respectively. In contrast, by leveraging well-pretrained ViT parameters, the visual register mechanism requires fewer than 3M samples for adaptation. The ablation results in Fig. 4 further demonstrate this advantages. 2) Addressing Visual Fragmentation. By leveraging the compact and comprehensive representation produced by visual registers, the proposed ReAtten module can effectively facilitate information exchange across sub-images. This is an important function that is notably lacking in Abstractor-like modules. Combined with ReCompact, it enables the visual register mechanism to simultaneously address the issues of visual redundancy and fragmentation. The ablation results in Tab. 5 validate the effectiveness of ReAtten.
@@
  • To address the issue of visual redundancy directly at the encoding phase, we introduce the Register-based Representation Compacting (ReCompact) mechanism. This mechanism enables vision encoder to capture rich visual information using only a minimal number of tokens, effectively eliminating redundancy and obviating the need for additional compression modules.
@@
  • We introduce FALCON, a model that leverages the proposed visual register technique. Comprehensive evaluations on high-resolution visual understanding benchmarks highlight the superiority of FALCON. Notably, FALCON achieves exceptional results with a remarkable reduction in visual tokens by 9×.
@@ 2.1 Multimodal Large Language Models
Research on MLLMs [30, 2, 1, 58, 59, 46] has made significant strides in recent years. By incorporating image inputs into pretrained LLMs [50, 51, 8], MLLMs demonstrate strong capabilities across various vision-language tasks. Typically, these models use a pretrained vision encoder [42, 61] to transform input images into a set of visual tokens. These tokens are then projected into the input space of the LLM through a projection module (e.g., MLP [32, 30], QFormer [9] and Abstractor [59]) and processed alongside text inputs. Building on the progress of MLLMs, recent studies have shifted from designing specialized methods for certain tasks [44, 45, 43] to employing MLLMs as generalist agents, capable of handling diverse practical tasks such as document analysis [56, 15, 63], online video assistant [24], game agents [27, 29], robot manipulation [21, 18, 39], and GUI agent [5, 3, 53]. Despite these advances, MLLMs still struggle with the fixed low resolution of pretrained image encoders, which can limit their effectiveness in handling high-resolution images.
@@ 3.2 Ensuring Visual Continuity via ReAtten
Figure 3: The progressive training pipeline. Stage 0: Align the frozen vision and language backbones by training only the MLP and registers. Stage 1: Tune all parameters with caption and OCR data on low-resolution for coarse alignment. Stage 2: Integrate high-resolution input along with fine-grained captions and OCR to achieve refined alignment. Stage 3: Froze ViT and introduce ReAtten module for high-resolution instruction tuning.
@@ 3.3 Learning Expressive Visual Representation via Progressive Training Pipeline
To empower the ReCompact mechanism with the capability to capture rich visual patterns and to enable smoother initialization of the ReAtten module under high-resolution conditions, we develop a specialized progressive training pipeline. As illustrated in Fig. 3, this pipeline progressively transitions from low to high resolution across four stages.
Stage 0: Static Alignment on Low-resolution. Since both the visual registers and MLP projector are randomly initialized, in stage 0, we aim to achieve a initial alignment between the pretrained vision encoder and the LLM. To facilitate this, we begin by using low-resolution images as inputs, which are directly resized to match the resolution of the vision encoder. We then statically pretrain the model by keeping both the vision encoder and LLM frozen, tuning only the visual registers and MLP using image caption data.
Stage 1: Coarse Alignment on Low-resolution. To better adapt the pre-trained vision encoder to our newly introduced ReCompact mechanism, in stage 1, we unfreeze all model parameters to enable a more thorough tuning. To facilitate the visual registers to progressively learn capturing visual information from coarse to fine granularity, we continue to use low-resolution inputs and employ high-quality, extended captions rich in visual content, along with full-text OCR data for training. This design allows the registers to acquire holistic scene understanding ability.
Stage 2: Refined Alignment on High-resolution. To further enhance the capacity of the visual register to capture fine-grained visual details, in stage 2, we transition the model to process high-resolution image for refined alignment. Additionally, we employ datasets that includes a series of fine-grained captions and OCR tasks, such as regional captions and text location. This facilitates the visual registers to capture comprehensive fine-grained details.
Stage 3: Instruction Tuning on High-resolution. To adapt the well-aligned visual register for a wide range of real-world tasks, in stage 3, we perform instruction tuning with a mixture of datasets covering various tasks, such as VQA, document understanding, and science reasoning. Additionally, the proposed ReAtten module is integrated to address visual fragmentation caused by cropping-based high-resolution image encoding. Since the vision encoder with ReCompact is well-aligned, the vision encoder is kept frozen and other parameters including ReAtten are trainable during instruction tuning. Notably, to facilitate a smoother initialization for the newly introduced ReAtten modules, we initialize parameters of each ReAtten module with the ViT self-attention parameters at the corresponding depth. This design enables ReAtten to effectively leverage pre-trained parameters, avoiding the need for extensive fine-tuning.
@@ 4 Experiments
We implement FALCON with SigLIP-L/16-384px [61] as the vision encoder and Llama-3.1-8B-Instruct [12] as the LLM backbone. A two-layer MLP module with GeLU [13] activation is served as the vision-language projection layer. The number of visual registers is set to 64, achieving remarkable visual token compression rates of 9×9$\times$9 ×. The maximum number of sub-images is set to 16.
@@ 4.1 Experimental Results
Table 1: Evaluation on MME-RealWorld Benchmark. “OCR”, “RS”, “DT”, “MO”, and “AD” stand for OCR in the Wild, Remote Sensing, Diagram and Table, Monitoring, and Autonomous Driving, respectively. “Avg” denotes the weighted average accuracy across all samples, while “Avg-C” denotes the weighted average accuracy across different domains. “$\text{Size}^{v}$” indicates the image resolution supported by the model, where ×$\times n$ signifies a cropping-based high-resolution processing approach with up to $n$ sub-images. “$\text{Token}^{c}$” denotes the number of visual tokens after compression. † For models utilizing hybrid vision encoders, “$\text{Size}^{v}$” represents the setting of encoder with the highest resolution. We gray out the models that are either closed-source or have at least twice the number of visual tokens compared to ours. Among these, the lighter gray denotes scores below ours. The underline indicates that the second performance is achieved. Notably, FALCON achieves state-of-the-art performance with just 64 tokens per sub-image, surpassing most models that have large amount of visual tokens and many advanced closed-source commercial models.
@@
Visual Tokens $\leq$ 1024 MiniGPT-v2 [4] $448^{2}$ $256$ 39.0 23.3 20.4 19.3 26.0 26.9 25.6 30.0 20.4 16.9 23.7 23.0 22.7
Qwen-VL-Chat [1] $448^{2}$ $256$ 32.4 15.1 15.6 22.1 15.1 20.8 20.1 28.6 13.6 16.5 24.6 22.0 20.8
@@
FALCON $384^{2}\times 16$ $\textbf{64}\times 16$ 66.4 49.6 47.2 40.4 36.5 50.3 48.0 55.4 39.2 44.4 34.5 40.7 43.4
Visual Tokens $\geq$2048 LLaVA-Next-Qwen72B [31] $336^{2}\times 4$ $576\times 4$ 37.1 29.1 27.7 29.4 18.0 29.0 28.3 17.2 34.2 27.3 29.7 27.9 27.1
@@
As shown in Tab. 1, FALCON significantly outperform other advanced MLLMs in both perception and reasoning tasks. FALCON achieves state-of-the-art performance with only 64 tokens per sub-image, resulting in a remarkable visual compression rate of 9 times. In addition, FALCON also outperforms most models that have twice as many visual tokens. Specifically, FALCON outperforms CogVLM2-llama3-Chat [14] in average scores on both perception and reasoning task, which uses 2304 tokens per-image. It also surpasses IXC2.5 [62] in average scores on the reasoning tasks, which has a maximun token counts of 9600. Notably, FALCON achieves higher average scores in both perception and reasoning tasks compared to three advanced commercial models, GPT-4o, GPT-4o-mini and Gemini-1.5-pro [49], with it notably outperforming all commercial models in the sub-tasks of remote sensing and monitoring. These results demonstrate that the ReCompact and ReAtten technique proposed by FALCON effectively eliminate visual redundancy and ensure visual continuity, allowing comprehensive and detailed visual information to be conveyed with a highly compact visual representation.
To further assess FALCON’s capabilities in high-resolution understanding, we conduct evaluations on more high-resolution benchmarks: V* [52], DocVQA [41], TextVQA [48] and ChartQA [40]. V* focuses on evaluating models’ capabilities for fine-grained attribute recognition and spatial reasoning. DocVQA challenges models to comprehend text-dense document images. ChartQA evaluates models’ proficiency in interpreting various types of chart, while TextVQA assesses the ability to recognize small text in real-world scenes. As shown in Tab. 2, FALCON demonstrates superior performance across all these benchmarks.
To evaluate the performance of FALCON across a broader range of domains, we further conduct experiments with diverse popular benchmarks. As shown in Tab. 4, POPE [25], ScienceQA [37], MMBench [34] and SEED-Bench [19] are selected to assess models’ capabilities in hallucination, scientific reasoning, and comprehensive ability. FALCON achieves SOTA performance across all these benchmarks. The results demonstrate that FALCON exhibits superior capabilities across a wide range of tasks.
Table 2: Evaluations on diverse high-resolution Benchmarks. “$\textbf{V}^{*}_{Avg}$” represents the weighted average score for the two sub-tasks in the $V^{*}$ benchmark. † denotes OCR inputs are utilized.
Model $\textbf{V}^{*}_{Avg}$ DocVQA TextVQA ChartQA
@@
Qwen-VL [1] - 65.1 63.8 59.0
InstructBLIP [9] 34.0 4.5 50.7 10.9
LLaVA-v1.5-7B [30] 48.7 8.5 58.2† 17.8
mPLUG-Owl2 [59] - 17.9 58.2† 22.8
@@
UReader [57] - 65.4 57.6 59.3
Monkey [28] - 66.5 65.8 59.0
LlaVA-FlexAttn [23] 54.4 - 48.9 -
S2-Wrapper-7B [47] 55.5 - 61.0 -
LlaVA-Next-Vicuna7B [31] 61.8 - 64.9 54.3
SliME-7B [66] - - 64.4 33.9
FALCON 64.4 74.4 66.5 68.8
Table 3: Ablation results on training data. We train FALCON with two-stage pipeline and exactly the same dataset as LLaVA-v1.5 [30]. The results confirm the superiority of the proposed method when the influence of training data is excluded.
Model SQA POPE MMB GQA TextVQA MM-Vet SEED-Img
LlaVA-v1.5-7B [30] 66.8 85.9 64.3 62.0 58.2 31.1 65.8
Ours 68.9 87.5 66.0 62.5 61.3 31.4 67.8